# thanks: Corresponding author

### v1

@@ IV Critical-exponent invariance
Critical exponents quantify the universal scaling near the black-hole critical point. They are conventionally defined with the horizon radius $r_{h}$, yet the interchangeability of $r_{\rm c}$ with $r_{h}$ suggests that one may equally define them using the circular-orbit radius $r_{\rm c}$. In this section we prove that, for both isobaric and isothermal constraints, the mapping $r_{\rm c}(r_{h})$ is locally monotonic at criticality, so the replacement $r_{h}\mapsto r_{\rm c}$ amounts to a smooth linear reparametrization in a neighborhood of the critical point. Consequently, critical exponent defined with $r_{\rm c}$ coincides with its counterpart defined with $r_{h}$.
The order-parameter exponent $\beta$ is defined from the coexistence gap of the horizon radius for $T<T_{c}$:
@@
with the mean-field value $\beta=1/2$. We will show that the circular-orbit radius inherits the same scaling, namely
$\Delta r_{c}\;:=\;r_{c}\!\big(r_{h,l}\big)-r_{c}\!\big(r_{h,s}\big)\;\sim\;|t|^{1/2}.$ (66)
We first perform a Taylor expansion of $P(r_{h},T)$ about the critical point $(r_{h,c},T_{c})$ up to the first nonvanishing orders allowed by the criticality conditions $\bigl.(\partial_{r_{h}}P)\bigr|_{c}=0$ and $\bigl.(\partial_{r_{h}}^{2}P)\bigr|_{c}=0$:
$P(r_{h},T)-P_{c}$ $=\left.(\partial_{T}P)\right|_{c}\,\Delta T+\left.(\partial_{T}\partial_{r_{h}}P)\right|_{c}\,\Delta T\,\Delta r$ (67)
$+\frac{1}{2}\left.(\partial_{r_{h}}^{2}\partial_{T}P)\right|_{c}\,\Delta T\,(\Delta r)^{2}$
$+\frac{1}{6}[\left.(\partial_{r_{h}}^{3}P)\right|_{c}+\left.(\partial_{T}\partial_{r_{h}}^{3}P)\right|_{c}\,\Delta T]\,(\Delta r)^{3}$
$+\mathcal{O}\!\big(\Delta T^{2},\,(\Delta r)^{4}\big),$
where $\Delta r:=r_{h}-r_{h,c}$ and $\Delta T:=T-T_{c}$. At fixed temperature $T$, differentiating with respect to $r_{h}$ gives
$\partial_{r_{h}}P(r_{h},T)$ $=\left.(\partial_{T}\partial_{r_{h}}P)\right|_{c}\,\Delta T$ (68)
$+\left.(\partial_{r_{h}}^{2}\partial_{T}P)\right|_{c}\,\Delta T\,(\Delta r)$
$+\frac{1}{2}[\left.(\partial_{r_{h}}^{3}P)\right|_{c}+\left.(\partial_{T}\partial_{r_{h}}^{3}P)\right|_{c}\,\Delta T]\,(\Delta r)^{2}$
$+\mathcal{O}\!\big(\Delta T^{2},\,(\Delta r)^{4}\big),$
On the coexistence line for $T<T_{c}$, the Eq. (65) gives
$\Delta r=(r_{h,l}-r_{h,c})-(r_{h,s}-r_{h,c})=\Delta r_{h}\sim|t|^{1/2},$ (69)
so each power counts as $\Delta r^{n}\sim|t|^{n/2}$. Substituting this into (67) and keeping the leading orders in $|t|$ yields
$\partial_{r_{h}}P(r_{h},T)$ (70)
$=\left.(\partial_{T}\partial_{r_{h}}P)\right|_{c}\,T_{c}\,t+\frac{1}{2}\left.(\partial_{r_{h}}^{3}P)\right|_{c}\,|t|+\mathcal{O}\!\big(|t|^{3/2}\big),$
Thus $\bigl(\partial P/\partial r_{h}\bigr)_{T}=\mathcal{O}(|t|)\to 0$ as $t\to 0^{-}$. Along an isotherm, our criterion
$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\!\left(\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{\epsilon}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{\epsilon}\right)$ (71)
combines with the first law
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T\,S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}}$ (72)
to give, as $t\to 0^{-}$,
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{c}\,S^{\prime}(r_{h,c})+\mathcal{O}(|t|),\qquad\frac{\mathrm{d}P}{\mathrm{d}r_{h}}=\mathcal{O}(|t|).$ (73)
Since $S^{\prime}(r_{h,c})>0$, $V>0$, and $\Gamma_{\epsilon}\neq 0$, $\partial_{M}\Phi_{\epsilon}$ and $\partial_{P}\Phi_{\epsilon}$ are finite at the critical point, we obtain the fixed-sign leading term
$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{T_{c}\,S^{\prime}(r_{h,c})}{\Gamma_{\epsilon}}\,\partial_{M}\Phi_{\epsilon}\ +\ \mathcal{O}(|t|),$ (74)
so $r_{c}(r_{h})$ is strictly monotonic for sufficiently small $|t|$. (In the photon case, $\partial_{P}\Phi_{0}=0$ further collapses the subleading $\mathcal{O}(|t|)$ contribution.)
Monotonicity (74) implies the first-order linearization at $r_{h,c}$,
$r_{\rm c}(r_{h})=r_{{\rm c},c}+C\,\bigl(r_{h}-r_{h,c}\bigr)+o\!\bigl(|r_{h}-r_{h,c}|\bigr),$ (75)
where $C:=\left.{dr_{c}}/{dr_{h}}\right|_{c}\neq 0$. Therefore on the coexistence line
$\Delta r_{c}=r_{c,l}-r_{c,s}=C\,\Delta r_{h}+o(\Delta r_{h})\sim|t|^{\beta},$ (76)
and the critical exponent is unchanged:
$\beta_{(r_{c})}=\beta_{(r_{h})}=\tfrac{1}{2}\quad\text{(mean-field)}.$ (77)
For the isobaric process, $r_{c}(r_{h})$ is globally monotonic, so the above conclusion applies verbatim. In summary, under either isobaric or isothermal constraints, the reparameterization $r_{h}\!\mapsto\!r_{c}$ does not change the critical exponents. This is consistent with the numerical findings of Wei et al. in the isobaric photon-sphere case [12], for which we have provided an analytic proof.
V Discussion
In this work, we derived the general criterion (21), applicable to any parameter path in spherically symmetric black holes, which clarifies how circular orbits encode thermodynamic behavior. The criterion reveals that the monotonic relation between the circular-orbit radius $r_{\mathrm{c}}$ and the horizon radius $r_{h}$ is influenced by two factors: the intrinsic properties of the black hole and its circular orbits, and the parameter path chosen in the extended thermodynamic space. While the former never breaks monotonicity, the latter plays a decisive role. Motivated by this, we investigated two thermodynamically significant paths–the isobaric and isothermal processes–using the $d$-dimensional RN-AdS black hole as a test case.
Our analysis demonstrates that the first law of thermodynamics directly constrains the monotonicity of $r_{\mathrm{c}}(r_{h})$. Under the isobaric constraint, the first law enforces $\mathrm{d}M/\mathrm{d}r_{h}\neq 0$, ensuring that $M(r_{h})$ and hence $r_{\mathrm{c}}(r_{h})$ remain monotonic. In contrast, under the isothermal constraint the first law allows $\mathrm{d}M/\mathrm{d}r_{h}=0$, which decouples the extrema of $M(r_{h})$ and $P(r_{h})$ so that a single mass $M$ corresponds to two distinct pressures $P$. This inevitably makes $r_{\mathrm{ps}}(r_{h})$ non-monotonic, since the photon-sphere radius $r_{\mathrm{ps}}$ is $P$-independent whereas $r_{h}$ depends on $P$. An exception occurs in the Sch-AdS limit ($q=0$), where $M(r_{h})$ remains monotonic; the emergence of electric charge $q$ thus breaks this monotonicity and, in turn, that of $r_{\mathrm{ps}}(r_{h})$.
Interestingly, for the ISCO case, both $r_{\mathrm{isco}}$ and $r_{h}$ depend explicitly on $P$, which might seem to favor monotonicity. However, for the RN-AdS family we find that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic for all $d$, including the Sch-AdS limit. This arises because the first law requires $\mathrm{d}M/\mathrm{d}r_{h}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ to share the same sign in certain domains, making the prefactor in the criterion (38) change sign and thus breaking monotonicity. Overall, our results indicate that $r_{\mathrm{c}}(r_{h})$ tends to be monotonic under isobaric evolution but non-monotonic under isothermal evolution.
It is worth noting that this non-monotonicity does not spoil the monotonic behavior of $r_{\rm c}(r_{h})$ in the local phase-transition region, nor does it affect its behavior near the critical point. As a result, the jump in $r_{\rm c}$ remains a reliable signature of the transition, and the critical exponents defined through $r_{\rm c}$ coincide with those obtained from $r_{h}$. These features are closely tied to the black hole first law and are expected to extend to more general classes of black holes.
These results fill an important gap in the study of the non-monotonic behavior of $r_{\mathrm{c}}(r_{h})$ and provide a more robust foundation for identifying black hole phase transitions through optical observables. They also offer an additional avenue for probing black hole thermodynamics: the monotonicity (or lack thereof) of $r_{\mathrm{c}}(r_{h})$ can be used to distinguish whether a black hole evolves along an isobaric or an isothermal process.
This idea is physically well motivated. By the entropy-increase law, the horizon radius $r_{h}$ must grow with time, and if $r_{\mathrm{c}}(r_{h})$ is non-monotonic, the resulting optical appearance of the black hole will exhibit correspondingly distinct behavior. Moreover, the influence of black hole thermodynamics on the monotonicity of $r_{\mathrm{c}}(r_{h})$ is itself a noteworthy feature deserving further investigation.
@@ Acknowledgements.
This work is supported in part by NSFC Grants No. 12165005 and No. 11961131013.
@@ Appendix A Preservation of Functional Behavior under Monotonic Reparameterization
$\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}=F^{\prime\prime}\bigl(x(y)\bigr)\,[x^{\prime}(y)]^{2}+F^{\prime}\bigl(x(y)\bigr)\,x^{\prime\prime}(y).$ (82)
At a stationary point $y_{m}$ where $F^{\prime}\bigl(x(y_{m})\bigr)=0$, this reduces to
$\left.\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}\right|_{y_{m}}=F^{\prime\prime}(x_{m})\,[x^{\prime}(y_{m})]^{2}.$ (83)
Since $x^{\prime}(y_{m})]^{2}>0$, it follows that
$F^{\prime\prime}(x_{m})>0$ $\ \Longrightarrow\ \widetilde{F}^{\prime\prime}(y_{m})>0,$ (84)
$F^{\prime\prime}(x_{m})<0$ $\ \Longrightarrow\ \widetilde{F}^{\prime\prime}(y_{m})<0.$
@@
If $F$ is identified with the thermodynamic temperature $T$ or pressure $P$ of a black hole, and the variables are smoothly and monotonically related as $x=r_{h}$ and $y=r_{c}$, then under the mapping $r_{h}\mapsto r_{c}$ the functions $T(r_{c})$ and $P(r_{c})$ retain their characteristic van der Waals–like $S$-shaped structure. As a result, during the small-to-large black-hole phase transition, the critical radius $r_{c}$ undergoes a finite discontinuous jump.
Appendix B Constructive counterexample in Sch-AdS demonstrating that $r_{\rm ps}(r_{h})$ can decrease
For the four-dimensional Sch-AdS black hole, the metric function reads

### v2

@@ IV Critical-exponent invariance
In the previous section, we showed that the mapping $r_{c}(r_{h})$ can exhibit global nonmonotonicity. On the one hand, this implies that encoding phase transitions in the circular orbit radius is not always reliable. On the other hand, from a thermodynamic viewpoint it raises a deeper issue: Whether such global nonmonotonicity can affect the universal critical behavior governed by the neighborhood of the critical point. Specifically, near criticality, the distinction between the coexisting phases becomes progressively weaker and vanishes exactly at the critical point. In conventional thermodynamics, such critical behavior is found to be universal and can be quantitatively characterized by critical exponents. For example, the classical Van der Waals system exhibits the order-parameter critical exponent $\beta=1/2$, which has also been shown to hold for many AdS black holes when the thermodynamic volume or the horizon radius is adopted to define the order parameter. Recently, Wei et al. demonstrated that in the isobaric ensemble the four-dimensional RN–AdS black hole yields the same value $\beta=1/2$ when the photon sphere radius is used to define the order parameter, and their numerical results for higher dimensional RN–AdS black holes remain very close to $1/2$ [19]. These findings motivate us to test whether such universal critical behavior persists in different ensembles, especially when the mapping $r_{c}(r_{h})$ becomes globally nonmonotonic. In this section we prove that, under both isobaric and isothermal constraints, the mapping $r_{c}(r_{h})$ is necessarily locally monotonic at criticality. As a result, the reparametrization $r_{h}\mapsto r_{c}$ reduces to a smooth linear transformation in a neighborhood of the critical point. This guarantees that the critical exponents defined using $r_{c}$ coincide exactly with those obtained from $r_{h}$, so the universal critical behavior is fully preserved.
The order-parameter critical exponent $\beta$ is defined from the coexistence gap of the horizon radius for $T<T_{c}$:
@@
We will show that the circular orbit radius inherits the same scaling, namely
$\Delta r_{c}\;:=\;r_{c}\big(r_{h,l}\big)-r_{c}\big(r_{h,s}\big)\;\sim\;|t|^{\beta}.$ (63)
The proof of Eq. (63) is straightforward. We first show that $r_{\rm c}(r_{h})$ is monotonic in the neighborhood of the critical point and then demonstrate that such a monotonic reparametrization leaves the critical exponents unchanged. Near the critical point one has ${\rm d}P/{\rm d}r_{h}\!\to\!0$, so that the isothermal criterion (38) reduces to
$\frac{{\rm d}r_{\rm c}}{{\rm d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\,\frac{{\rm d}M}{{\rm d}r_{h}}\,\partial_{M}\Phi_{\epsilon},$ (64)
which coincides exactly with the isobaric criterion (27). Consequently, in both ensembles the monotonic behavior of $r_{\rm c}(r_{h})$ near the critical point is entirely determined by that of $M(r_{h})$. In the isobaric ensemble, the first law enforces ${\rm d}M/{\rm d}r_{h}>0$, and $r_{\rm c}(r_{h})$ remains monotonic. In the isothermal ensemble, the first law gives
$\frac{{\rm d}M}{{\rm d}r_{h}}=T_{0}\,S^{\prime}(r_{h})+V\,\frac{{\rm d}P}{{\rm d}r_{h}},$ (65)
which simplifies near the critical point to
$\frac{{\rm d}M}{{\rm d}r_{h}}=T_{0}\,S^{\prime}(r_{h}),$ (66)
again ensuring ${\rm d}M/{\rm d}r_{h}>0$ and hence the monotonicity of $r_{\rm c}(r_{h})$. Therefore, in both the isobaric and isothermal ensembles,
$C:=\left.\frac{{\rm d}r_{\rm c}}{{\rm d}r_{h}}\right|_{c}\neq 0.$ (67)
We next expand $r_{\rm c}(r_{h})$ about the critical radius $r_{h,c}$ as
$r_{\rm c}(r_{h})=r_{{\rm c},c}+C\,\bigl(r_{h}-r_{h,c}\bigr)+o\!\left(|r_{h}-r_{h,c}|\right).$ (68)
This immediately yields
$\Delta r_{\rm c}=r_{c,l}-r_{c,s}=C\,\Delta r_{h}+o(\Delta r_{h})\sim|t|^{\beta},$ (69)
which shows that, provided $C\neq 0$, the leading behavior of $\Delta r_{\rm c}$ near the critical point is governed by $\Delta r_{h}$. As a result, the critical exponent remains unchanged:
$\beta_{(r_{c})}=\beta_{(r_{h})}.$ (70)
This result is noteworthy: Although $r_{\rm c}(r_{h})$ may exhibit global nonmonotonic behavior, it must be monotonic in the neighborhood of the critical point, which ensures that the critical exponent remains invariant under reparametrization. This shows that the circular orbit radius encodes not only the existence of phase transitions but also their universal critical properties. Moreover, it is worth emphasizing that this proof does not rely on the specific form of the metric (23), but only on the criterion (38) and the first law. The result therefore applies, in principle, to any black hole with $\partial_{M}\Phi_{\epsilon}\neq 0$, which points to a more fundamental connection between black hole gravitational geometry and thermodynamics.
V Summary and Discussion
In this paper, we derived the monotonicity criterion (21) that applies to arbitrary thermodynamic paths in the parameter space of any static, spherically symmetric black hole, thereby clarifying when the mapping between the horizon radius $r_{h}$ and the circular orbit radius $r_{\rm c}$ is monotonic. In this framework, the monotonicity of $r_{\rm c}(r_{h})$ is governed by two factors: (i) the intrinsic geometric/orbital properties of the black hole and (ii) the chosen path in the extended thermodynamic parameter space. For any nonextremal black hole admitting an unstable photon sphere or an ISCO, the first factor does not by itself induce nonmonotonicity, whereas the second is decisive. Motivated by this, we take the dd–dimensional RN–AdS black hole as a concrete example and analyze two thermodynamically cases: the isobaric and isothermal ensembles.
Our results show that the monotonicity of $r_{\rm c}(r_{h})$ is completely different in the two ensembles, and this contrast is closely tied to the first law of black hole thermodynamics. In the isobaric ensemble, the sign of $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ is determined by $\partial_{M}\Phi_{\epsilon}$ together with $\mathrm{d}M/\mathrm{d}r_{h}$. For RN–AdS, $\partial_{M}\Phi_{\epsilon}\neq 0$, while the first law enforces ${\rm d}M/{\rm d}r_{h}>0$, so ${\rm d}r_{\rm c}/{\rm d}r_{h}$ does not change sign and $r_{\rm c}(r_{h})$ is globally monotonic (both $r_{\rm ps}(r_{h})$ and $r_{\rm isco}(r_{h})$ are strictly increasing). This monotonic increase guarantees that, whenever a phase transition occurs, $r_{\rm c}$ jumps together with $r_{h}$, thereby ensuring that the circular orbit radius reliably encodes the phase transition information. By the universality of the first law, this conclusion is not restricted to RN–AdS but extends to any black hole with $\partial_{M}\Phi_{\epsilon}\neq 0$.
In contrast, in the isothermal ensemble the sign of $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ is controlled by four factors: $\partial_{M}\Phi_{\epsilon}$, $\partial_{P}\Phi_{\epsilon}$, $\mathrm{d}M/\mathrm{d}r_{h}$, and $\mathrm{d}P/\mathrm{d}r_{h}$. For the photon sphere one has $\partial_{P}\Phi_{0}=0$, so the sign of $\mathrm{d}r_{\rm ps}/\mathrm{d}r_{h}$ again depends only on $\partial_{M}\Phi_{0}$ and $\mathrm{d}M/\mathrm{d}r_{h}$, but now the first law allows $\mathrm{d}M/\mathrm{d}r_{h}$ to cross zero. As a result, $\mathrm{d}r_{\rm ps}/\mathrm{d}r_{h}$ may change sign and $r_{\rm ps}(r_{h})$ can become nonmonotonic. This is precisely what happens for the RN–AdS black hole, whose $r_{\rm ps}(r_{h})$ indeed exhibits nonmonotonic behavior, whereas the Sch–AdS case is exceptional: The absence of charge eliminates any zero of $\mathrm{d}M/\mathrm{d}r_{h}$, so $r_{\rm ps}(r_{h})$ remains monotonic. For the ISCO, one has $\partial_{P}\Phi_{1}>0$, so the sign of $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$ is jointly controlled by $\mathrm{d}M/\mathrm{d}r_{h}$ and $\mathrm{d}P/\mathrm{d}r_{h}$. The first law forces these two derivatives to share the same sign in certain ranges of $r_{h}$, thereby making sign changes in $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$ possible. This possibility is realized explicitly in RN–AdS and Sch–AdS, where $r_{\rm isco}(r_{h})$ is fully nonmonotonic. Given the universality of the first law, we expect similar nonmonotonic behavior to appear for other black holes evolving along isotherms. In principle, such nonmonotonicity could reduce the jump in $r_{\rm c}$ at the phase transition, or even eliminate it, thereby undermining the reliability of using $r_{\rm c}$ to encode the phase transition.
Fortunately, we have shown that, for RN–AdS, the phase transition always takes place on an increasing branch of $r_{\rm c}(r_{h})$, so the reliability of using $r_{\rm c}$ to encode the phase transition is not compromised. Whether this remains true for other black holes requires further case-by-case checks: One must first establish whether $r_{\rm c}(r_{h})$ becomes nonmonotonic in the isothermal ensemble and then determine whether this nonmonotonicity alters the jump amplitude $\Delta r_{\rm c}=r_{{\rm c},l}-r_{{\rm c},s}$ across the phase transition. Both steps rest on the monotonicity criterion (21), so in this sense our work provides an effective theoretical framework for assessing the observability of black hole phase transitions.
It is also worth emphasizing that the nonmonotonic behavior of $r_{\rm c}(r_{h})$ in the isothermal ensemble has, to our knowledge, gone unnoticed so far. Thus, previous studies have largely focused on extracting phase transition information from black hole optical appearances, whereas we now identify an additional channel: The nonmonotonic evolution of the optical appearance can provide a diagnostic of the underlying thermodynamic ensemble. Since any nonmonotonicity of $r_{\rm c}(r_{h})$ should be reflected in the black hole image, such a test is feasible in principle. Of course, from an observational perspective, the relevant observables are the shadow and the accretion–disk image rather than $r_{\rm ps}$ or $r_{\rm isco}$ directly. Thus our next step is to analyze how the shadow and accretion–disk images evolve under isobaric versus isothermal processes.
Finally, we prove that whenever a black hole exhibits critical behavior, the first law alone guarantees invariance of the critical exponents under the reparametrization $r_{h}\!\to r_{\rm c}$ for any spacetime satisfying $\partial_{M}\Phi_{\epsilon}\neq 0$. This shows that circular particle orbits not only encode phase transitions but also retain the universal critical behavior of black hole thermodynamics. This generality hints at a deeper geometric–thermodynamic correspondence that merits further study.
@@ Acknowledgements.
We would like to thank Zhongzhinan Dong and Jiawei Chen for helpful discussions. This work is supported in part by NSFC Grants No. 12165005 and No. 11961131013.
@@ Appendix A Preservation of Functional Behavior under Monotonic Reparameterization
$\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}=\frac{\mathrm{d}^{2}F}{\mathrm{d}x^{2}}\left(\frac{\mathrm{d}x}{\mathrm{d}y}\right)^{2}+\frac{\mathrm{d}F}{\mathrm{d}x}\frac{\mathrm{d}^{2}x}{\mathrm{d}y^{2}}.$ (75)
At a stationary point $y_{m}$ where ${\mathrm{d}F}/{\mathrm{d}x}=0$, this reduces to
$\left.\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}\right|_{y_{m}}=\left.\frac{\mathrm{d}^{2}F}{\mathrm{d}x^{2}}\right|_{x_{m}}\left.\left(\frac{\mathrm{d}x}{\mathrm{d}y}\right)^{2}\right|_{y_{m}}.$ (76)
Since $\left({\mathrm{d}x}/{\mathrm{d}y}\right)^{2}>0$, it follows that
$\left.\frac{\mathrm{d}^{2}F}{\mathrm{d}x^{2}}\right|_{x_{m}}>0$ $\ \Longrightarrow\ \left.\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}\right|_{y_{m}}>0,$ (77)
$\left.\frac{\mathrm{d}^{2}F}{\mathrm{d}x^{2}}\right|_{x_{m}}<0$ $\ \Longrightarrow\ \left.\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}\right|_{y_{m}}<0.$
@@
If $F$ is identified with the thermodynamic temperature $T$ or pressure $P$ of a black hole, and the variables are smoothly and monotonically related as $x=r_{h}$ and $y=r_{c}$, then under the mapping $r_{h}\mapsto r_{c}$ the functions $T(r_{c})$ and $P(r_{c})$ retain their characteristic van der Waals–like $S$-shaped structure. As a result, during the small-to-large black hole phase transition, the critical radius $r_{c}$ undergoes a finite discontinuous jump.
Appendix B Constructive counterexample in Sch–AdS demonstrating that $r_{\rm ps}(r_{h})$ can decrease
For the four-dimensional Sch–AdS black hole, the metric function reads